OpenStudy (anonymous):

Let R be the region bounded by the graphs of y=cos((pi x)/2) and y=x^2-(26/5)x+1.

OpenStudy (anonymous):

Let R be the region bounded by the graphs of y=cos((pi x)/2) and y=x^2-(26/5)x+1. A. Find the area of R. B. The vertical line x=k splits the region R into two equal parts. Write, but do not solve, an equation involving integrals that solves for k. C. The region R is the base of a solid. For this solid, each cross section perpendicular to the x-axis is a square. Find the volume of the solid. D. The region R models the surface of a piece of glass. At all points R at a distance x from the y-axis, the thickness of the glass is given by h(x)=5-x. Find the volume of the glass. http://goo.gl/vSltp